Hello out there,I need some help with some problems,the first one is,A right triangle has legs of lengths 10 and 24.What is the length of the hypotenuse?Round to the nearest tenths.The next one is,A right triangle has a hypotenuse equal to 35 and one leg of length 30.Find the length of the other leg to the nearest tenth.The next one is,Write -4 times y times y times y using exponents.And the last one is,Write 3cd(to the third power) as a multiplication expression.Thanks to whoever can help.
Answer by checkley75(3666)

SIDES OF A TRIANGLE IS A^2+B^2=C^2 C BEING THE HYPOTENUSE.
10^2+24^2=C^2
100+576=C^2
C^2=676
C=SQRT676
C=26 IS THE LENGTH OF THE HYPOTENUSE.
----------------------------------------------------------------
35^2=30^2+X^2
1225=600+X^2
X^2=1225-600
X^2=625
X=SQRT625
X=25 IS THE LENGTH OF THE SECOND SIDE.
--------------------------------------------------------------
-4*Y*Y*Y=-4Y^3
--------------------------------------------------------------
3CD*3CD*3CD=27C^3D^3